OpenStudy (anonymous):

can anyone help me with this question? differentiate the implicit equation. 60x =15y + 0.0225*x^2* y

terenzreignz (terenzreignz):

Hey :) Whenever you're differentiating implicitly, with respect to x, you differentiate the y's as if they were x's, but when you're done, multiply it with dy/dx

terenzreignz (terenzreignz):

For instance, the 60x its derivative is just 60, so that's that for the left-hand side in 15y, its derivative, is 15 but since it was a y, you multiply it with dy/dx

OpenStudy (anonymous):

so for 0.0225*x^2*y, is the derivative (0.0225*x^2*dy/dx+0.0225*2x*y)?

terenzreignz (terenzreignz):

Yeah, actually :D

OpenStudy (anonymous):

Oh, Thanks! there's another part of this question which requires to substitute x=20 & y=50 into the dy/dx & the given answer is 1.875 but somehow i could only get 0.625

terenzreignz (terenzreignz):

hmm? I'm afraid I don't quite get it.

terenzreignz (terenzreignz):

okay, let's try it... \[\large 60 = 15\frac{dy}{dx}+0.0225x^2\frac{dy}{dx}+0.045xy\]

terenzreignz (terenzreignz):

Now, plug in x=20 and y=50 \[\large 60 = 15\frac{dy}{dx}+0.0225(20)^2\frac{dy}{dx}+0.045(20)(50)\]

terenzreignz (terenzreignz):

\[\large 60 = 15\frac{dy}{dx}+9\frac{dy}{dx}+45\]

terenzreignz (terenzreignz):

Did I miss anything?

terenzreignz (terenzreignz):

\[\large 60-45 = \frac{dy}{dx}(15+9)\]

terenzreignz (terenzreignz):

I'm getting 0.625 too That's strange... Maybe you're just too smart? :)

OpenStudy (anonymous):

really thanks for your help! I just know that the model answer was incorrect. XP

terenzreignz (terenzreignz):

Haha right